OpenStudy (anonymous):

7 3/10 + 6 1/3+ 2 7/10=

OpenStudy (lizz123):

The answer is 16 1/3

sammixboo (sammixboo):

Let's start with adding 7 3/10 and 2 7/10 together since they already have the same denominator. To make this easier we most make 7 3/10 and 2 7/10 imroper fractions. 7 3/10 = 73/10 2 7/10 = 27/10 Now add the numerators together